OpenStudy (anonymous):

Can someone please help me how to do this problem Algebra 2 xx Medal awarded! 1. A ball rebounds to a height of 30.0 cm on the third bounce and to a height of 5.2 cm on the sixth bounce. a. Find two points and plug into the point-ratio form y=y1*b^x-x1, using r for the ratio. Let x represent the bounce number, and let y represent the rebound height in centimeters. b. Solve the equation for b. Plug b and one of the two points into the point*ratio form to get an equation c.what height was the ball dropped from?

OpenStudy (anonymous):

i am sure we can do this, but what the monkey is the "point - ratio" formula?

OpenStudy (anonymous):

the point ratio formula is y=y1*b^x-x1

OpenStudy (anonymous):

ok i see i think it is \[\large y=y_1b^{x-x_1}\] right?

OpenStudy (anonymous):

yes yes :)

OpenStudy (anonymous):

\[y=30b^{x-3}\] and \[y=5.2b^{x-6}\] like that right?

OpenStudy (anonymous):

lets see if we can solve for \(b\)

OpenStudy (anonymous):

well one point is (x1,y1) and the other point goes in as (x,y)

OpenStudy (anonymous):

I think the points are (30,3) and (5.2, 6)

OpenStudy (anonymous):

\[30b^{x-3}=5.2b^{x-6}\] divide and get \[b^{x-3}=\frac{5.2b^{x-6}}{30}\] divide by \(b^{x-6}\) and get \[\frac{b^{x-3}}{b^{x-6}}=\frac{5.2}{30}\]

OpenStudy (anonymous):

you have the x and y backwards

OpenStudy (anonymous):

no matter, however, we are almost done

OpenStudy (anonymous):

so the two points would have to be (3,30) and (6,5.2)?

OpenStudy (anonymous):

\[\large\frac{b^{x-3}}{b^{x-6}}=b^{x-3-(x-6)}=b^3\]

OpenStudy (anonymous):

yes, the number of bounces is the \(x\) and the height is the \(y\)

OpenStudy (anonymous):

so finally we have \[\large b^3=\frac{5.2}{30}\] making \[\large b=\sqrt[3]{\frac{5.2}{30}}\] whatever that is

OpenStudy (anonymous):

i am sure there is an easier way to do this, but no matter

OpenStudy (anonymous):

now you need a calculator to find \(b\)

OpenStudy (anonymous):

looks like \(b\) is about \(0.5576\) http://www.wolframalpha.com/input/?i=cubed+root+%285.2%2F30%29

OpenStudy (anonymous):

@Mertsj you ever seen anything like this before?

OpenStudy (anonymous):

oh I see.. since c is asking the height that the ball was dropped from would i have to plug that value in or..

OpenStudy (anonymous):

ok so we have the answer to a and b right?

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

now you can use \[y=30\times (.5576)^{x-3}\]

OpenStudy (anonymous):

if you want the height the ball was dropped from, that is the "zeroth" bounce or \(x=0\) plug (0\) in to \[y=30\times (.5576)^{x-3}\] get \[y=30\times (.5576)^{-3}\] then a calculator

OpenStudy (anonymous):

i have to say this is a first for me

OpenStudy (anonymous):

oh wow .. :) you are very smart though:-)) it is kinda of confusing tbh

OpenStudy (mertsj):

@satellite73 You're too smart for your own good.

OpenStudy (anonymous):

lol hardly you ever seen this before?

OpenStudy (anonymous):

usually you would say (i think) \(ar^3=30\) and \(ar^6=5.2\) then solve

OpenStudy (anonymous):

solve for ar?

OpenStudy (anonymous):

solve for \(a\) and \(r\) but not matter, we are done

OpenStudy (anonymous):

*no matter

OpenStudy (anonymous):

this method is much easier i think \[\frac{ar^6}{ar^3}=\frac{5.2}{30}\] or \[r^3=\frac{5.2}{30}\] so \[r=\sqrt[3]{\frac{5.2}{30}}\] same as before exactly but i used \(r\) instead of \(b\) and solved quicker

OpenStudy (anonymous):

Oh i see.. so just to clarify :) for c it would be ar^3=30 and ar^b=5.2?

OpenStudy (anonymous):

forget what i just wrote we already solved for \(b\) and we get c via \[y=30\times (.5576)^{-3}\]

OpenStudy (anonymous):

and a calculator

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=30%28.5576%29^%28-3%29 looks like about 173

OpenStudy (anonymous):

ok thanks again so much you've been really very helpful! :)) xx